OpenStudy (anonymous):

Which statement is true about the difference square root of 7 - square root of 28? It is rational and equal to -2. It is rational and equal to 0. It is irrational and equal to -2square root of 7. It is irrational and equal tosquare root of 7.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@heyyeahyou

OpenStudy (anonymous):

@precious13

OpenStudy (mathmale):

It'd be best, I think, to reduce Sqrt(28) by first re-writing it as Sqrt(4)*Sqrt(7). Please simplify this.

OpenStudy (anonymous):

i dont get it @mathmale

OpenStudy (ranga):

Are you sure the last choice does not have a minus sign missing?

OpenStudy (anonymous):

yeah im sure @ranga

OpenStudy (ranga):

Is the question: \[\Large \sqrt{7} - \sqrt{28}\]

OpenStudy (anonymous):

\[Which statement is true about the difference 2\sqrt{7} -\sqrt{28}\]

OpenStudy (ranga):

that is different from the original posting. \[2\sqrt{7} - \sqrt{28} = 2\sqrt{7} - \sqrt{4*7} = 2\sqrt{7} - 2\sqrt{7} = 0\]

OpenStudy (anonymous):

It is rational and equal to -2. It is rational and equal to 0. It is irrational and equal to -2square root of 7. It is irrational and equal tosquare root of 7.

OpenStudy (anonymous):

wich one is it @ranga

OpenStudy (ranga):

The answer is zero. which one says "equal to 0"?

OpenStudy (anonymous):

thank you so much @ranga can you help me on one more

OpenStudy (ranga):

I can assist but I am not supposed to be solving it.

OpenStudy (anonymous):

yeah i know im doing the work

OpenStudy (anonymous):

Simplify: square root of 3 times square root of 21 square root of 24 square root of 63 3 square root of 7 None of the above

OpenStudy (anonymous):

i think this problem is confusing

OpenStudy (ranga):

I will give a general form and you can use that solve this problem: \[\Large \sqrt{p} * \sqrt{q}= \sqrt{p*q}\]

OpenStudy (anonymous):

hi